Studies indicate that the price elasticity of demand for cigarettes is 0.2. If a pack of cigarettes currently costs $5.00. and the government wants to reduce smoking by 40% by how much more should the government tax cigarettes?


1
Expert's answer
2021-03-17T18:32:34-0400
"p_1=5"

"p_2=5+t"

"\varDelta p=\frac {p_2-p_1}{p_1}=\frac{t}{5}"


"E=\frac{\varDelta Q}{\varDelta p}"

"t=15"